OpenStudy (anonymous):

3(x - 2) + 4(2x - 6) = 6(x - 4) + 8(2x + 1) -7/3?

OpenStudy (aroub):

please odubo try to do them

OpenStudy (anonymous):

i did my answer is below the problem

OpenStudy (aroub):

ok so i didnt see the answer anyone help him if its right or wrong i have to go bye good night

OpenStudy (anonymous):

3x - 6 + 8x - 24 = 6x -24 + 16x + 8 -11x = 14 x =- 14/11

OpenStudy (anonymous):

thats wrong

OpenStudy (aroub):

so its wrong